Circular No. 6795

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997eh IN ANONYMOUS GALAXY J. Maza, Department of Astronomy, University of Chile, reports the

discovery of a supernova (B ∼ 18.5) by Roberto Antezana on a plate taken by L. Gonzalez using the Maksutov telescope at Cerro El Roble on Nov. 25.246 UT. SN 1997eh is located at α = 5^{h}49^{m}04^{s}.59, δ = −24^{o}29′01′′.6 (equinox 2000.0), which is 6′′.4 west and 0′′.8 north of the nucleus of its host galaxy. Confirming images of the supernova were obtained by E. Costa (University of Chile) via a CCD frame taken with the Cerro Tololo (CTIO) 1.5-m telescope and by Gonzalez at Cerro El Roble, both on Nov. 27.25.

R. Covarrubias and N. Suntzeff, CTIO, report that spectrophotometry (range 350–700 nm, resolution 0.3 nm), obtained by R. C. Smith on Dec. 23.3 UT with the CTIO 1.5-m telescope, shows SN 1997eh to be a type-Ia supernova, ∼ 6 weeks after maximum. A very noisy spectrum of the galaxy suggests a possible redhift of z ∼ 0.04.

SUPERNOVA 1997ei IN NGC 3963 S. Nakano, Sumoto, Japan, reports a discovery by M. Aoki, Tsukioka-

cho, Toyama, of an apparent supernova (mag 16.5) on six frames taken on Dec. 23 (limiting mag ∼ 18.0) with a 0.43-m f/6 reflector (+ unfiltered CCD) and confirmed on three frames taken on Dec. 24. SN 1997ei is located at α = 11^{h}54^{m}59^{s}.98, δ = +58^{o}29′26′′.4 (equinox 2000.0), which is 10′′ east and 10′′ south of the center of NGC 3963. Eleven previous frames taken by Aoki from 1996 Dec. 21 to 1997 Nov. 3 do not show an object at this position, though a possible faint image is present on a frame taken on 1997 Dec. 5; nothing is present at this position on the Digital Sky Survey.

GRB 971214 M. Boer, Centre d’Etude Spatiale des Rayonnements, CNRS, Toulouse;

B. A. Roberts and R. Malina, Center for EUVE Astrophysics, Berkeley; M. Feroci and L. Piro, Istituto di Astrofisica Spaziale, CNR, Rome; and K. Hurley, Space Sciences Laboratory, University of Berkeley, communicate: “EUVE observed the position of the x-ray counterpart of GRB 971214 between Dec. 15.8 and 16.16 UT. No source was detected during the 2000-s observation made with the Lexan (10.0 nm) filter. The 3σ upper limit to the countrate is 0.002 counts/s. Assuming a hydrogen column density of 10^{20} cm^{−2}, this corresponds to an upper limit to the flux of ∼ 1.7 × 10^{−13} erg cm^{−2} s^{−1}.”

1997 December 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green